Suppose the sequence x(n) = 5 * n-15, n-0,-. . , 16 is the input to the filter with the impulse response h(n) = (2,5,1). (a) Compute the output y/i(n) using MATLAB. ] help (numbers). b) Use the overlap-save method to compute the filter output using MATLAB. Suppose L 5 new data points are employed in each block. Use MATLAB function ovrlpsav ] help (numbers). What is the filter size? help (numbers). What is the block size (the size of the DFT)? help (numbers). help (numbers) How many symbols are discarded in each block due to aliasing? How many symbols of the old record are saved and become the first symbols of the new record? Compare with your result in (a). c) Suppose overlap-add method is employed with L What is the input data block size? help (numbers). help (numbers). What is the size of the DFT? help (numbers)
Answer #1

x(n)=[ -15 , -10 , -5 , 0 , 5 , 10 , 15 , 20 , 25 , 30 , 35 , 40 , 45 , 50 , 55 , 60 , 65 ];

h(n)=[ 2 5 1];

(a) y1(n) = x(n) * h(n) = [ -30 , -95 , -75 , -35 , 5 , 45 , 85 , 125 , 165 , 205 , 245 , 285 , 325 , 365 , 405 , 445 , 485 ,385 ,65]

(b) yos(n) = [ -30 , -95 , -75 , -35 , 5 , 45 , 85 , 125 , 165 , 205 , 245 , 285 , 325 , 365 , 405 , 445 , 485 ,385 ,65]

filter size is 17+3-1=19.

block size (N)=L+M-1=5+3-1=7.

M-1=3-1=2 symbols are discarded in each block due to aliasing.

N-(discarded symbols)=7-2=5 symbols of old block are saved and becomes the first symbols of new record.

(c) input data block size is 5.

size of DFT = L+M-1=5+3-1=7.